i appreciate all your help! thank you so much
At least one of the answers above is NOT correct. (1 point) How much money do winners go home with from the television quiz s 1

(1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 150 A.C. A sample of 5 Egyptian skull breadt2
Answer #1

I have solved incorrect part i.e. part b.

The provided sample mean is à = 134.8 and the sample standard deviation is s = 1.7889. The size of the sample is n = 5 and thIt should be

\large 134.8-2.776\frac{1.7889}{\sqrt{5}}\: \, \: to\: \:\: 134.8+2.776\frac{1.7889}{\sqrt{5}}

Dear Students,
I am waiting for your feedback. I have given my 100% to solve your problem. If you satisfied with my answer then please please hit like.
Thank You